OpenStudy (anonymous):

how do you integrate e^(-sqrt(x))/(4 sqrt(x)) dx

OpenStudy (amistre64):

very carefully :)

OpenStudy (amistre64):

e^(-sqrt(x)) ---------- dx eh.... can we make a u sub and say... 4sqrt(x) u = sqrt(x) ?

OpenStudy (anonymous):

yes i used u = -sqrt(x)

OpenStudy (amistre64):

du = 1/2sqrt(x) dx dx = du 2sqrt(x)

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

2sqrt(x) e^-u du --------------- ....murmur 4u

OpenStudy (anonymous):

now factor out 1/2?

OpenStudy (amistre64):

still have that sqrt(x) hanging in there causing troubles

OpenStudy (amistre64):

or can we simply equate that to a -u?

OpenStudy (amistre64):

e^-u du ----------- then right? 2

OpenStudy (anonymous):

wait what?

OpenStudy (amistre64):

2sqrt(x) = 2u right? since u = sqrt(x)

OpenStudy (anonymous):

right

OpenStudy (amistre64):

2u 1 -- = -- if its doable 4u 2

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

then all we are missing is a -1 to integrate with.... so I propose this set up: -(1/2) [S] -e^-u du

OpenStudy (amistre64):

which integrates to: -e^-u ------ +C lets check that thought and see if it works 2

OpenStudy (amistre64):

-e^-sqrt(x) --------- derives to what? 2

OpenStudy (amistre64):

- 1 -e^-sqrt(x) ------- * -------------- looks good to me 2sqrt(x) 4

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

i got lucky lol ;) youre welcome

OpenStudy (anonymous):

I just did this problem yesterday for calc, I was going to help but Amitre is far better at explaining then I am.

OpenStudy (anonymous):

:) hey bobby

OpenStudy (amistre64):

:) i just winged it; threw what i had at it to see what stuck ;)

OpenStudy (anonymous):

Woah who is this??

OpenStudy (anonymous):

lols chelsea

OpenStudy (anonymous):

fādāi means day dream in chinese